Amos and Micah are brothers. Amos is two years less than twice Mica's age. The sum of there age is 16. how old are they
Answer by ReadingBoosters(3246)

Set variables
x = amos
y = mica
...
Given
x = 2y - 2
x+y=16
...
Substitute x into the second equation
...
2y - 2 + y = 16
3y - 2 = 16
3y = 16 + 2
3y = 18
y = 6 is Mica's age
x = 2(6)-2 = 10 is Amos' age
...
Check
6 + 10 = 16
.........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.
www.MyHomeworkAnswers.com